[figure: 9.7 cm]

QUANTUM PHASE TRANSITIONS IN FRUSTRATED FERRO-

AND FERRIMAGNETS

J. Richter^{a}, N.B. Ivanov^{b} and U. Schollwöck^{c} aInstitute for Theoretical Physics, University Magdeburg, Germany bInstititute for Solid State Physics, Sofia, Bulgaria cSektion Physik, LMU München, Germany

Low-dimensional frustrated quantum magnets have attracted much attention in recent time [see, e.g., 1,2] since they are candidates for zero-temperature quantum orderdisorder transitions driven by frustration. We study the ground state properties of frustrated Heisenberg ferro- and ferrimagnets using density matrix renormalization group and exact diagonalization techniques. The frustration is introduced by competing second neighbor antiferromagnetic bonds J_{2}. In the limit of weak frustration for both systems the ground state belongs to macroscopic total spin (S ∝ N) and exhibits magnetic long-range order. On the other side, the ground state is a singlet (S = 0) in the limit of large J_{2}. We find indications that this transition to a singlet state is of first order in the frustrated ferromagnet in D = 1, D = 2 and D = ∞. On the other hand, in the frustrated ferrimagnet our results indicate also a first order transition in D = 1 but a second order transition in higher dimensions.

[1] D.V. Dimitriev, V.A. Krivnov and A.A. Ovchinnikov, Phys. Rev. B 56, 5985 (1997)

[2] N.B. Ivanov, J. Richter and U. Schollwöck, Phys.Rev. B 58, 14456 (1998)

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : oral

Corresponding author : J. Richter

Address for correspondence : Institute of Theoretical Physics University Magdeburg Postfach 4120 D-39016 Magdeburg Germany

Email address : johannes.richter@physik.uni-magdeburg.de